OpenStudy (anonymous):

Rebecca leaves on a long trip driving at a steady rate of 45 mph. Her sister Allison leaves from the same location traveling to the same destonation 2 hours later. She drives at a steady rate of 60 mph. How long after Allison leaves home will she catch up to Rebecca

OpenStudy (anonymous):

@dan815 @dtan5457 @dejmccray

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

@dtan5457 could you help me on these questions i have? Please i really need you to

OpenStudy (dtan5457):

45x=60(x-2) 45x=60x-120 -15x=-120 solve for x

OpenStudy (anonymous):

A. 9 Hours B.8 Hours C. 7 Hours D.6Hourss

OpenStudy (dtan5457):

I see the answer, just solve for x, lol

OpenStudy (anonymous):

8

OpenStudy (dtan5457):

yes